Find the equation of the line that contains (0,2) and (-4,5)
Answer by mananth(16946)

x1 y1 x2 y2
0 * 2 -4 5

slope m = (y2-y1)/(x2-x1)
( 5 - 2 )/( -4 - 0 )
( 3 / -4 )
m= - 3/ 4

Plug value of the slope and point ( 0 , 2 ) in
Y = m x + b
2.00 = -0 + b
b= 2 - -0
b= 2
So the equation will be
Y = - 3/4 x + 2
m.ananth@hotmail.ca